### v4

@@ 1 Introduction
Ordinal data are ubiquitous in psychology and related fields. With such data, e.g., arising from responses to rating scales, it is often recommended to estimate correlation matrices through polychoric correlation coefficients (e.g., Foldnes & Grønneberg,, 2022; Garrido et al.,, 2013; Holgado-Tello et al.,, 2010). The resulting polychoric correlation matrix is an important building block in subsequent multivariate models like factor analysis models and structural equation models (SEMs), as well as in explanatory methods like principal component analysis, multidimensional scaling, and clustering techniques (see, e.g., Mair,, 2018, for an overview). An individual polychoric correlation coefficient is the population correlation between two underlying latent variables that are postulated to have generated the observed categorical data through an unobserved discretization process. Traditionally, it is assumed that the two latent variables are standard bivariate normally distributed (Pearson,, 1901) to estimate the polychoric correlation coefficient from observed ordinal data. Estimation of this latent normality model, called the polychoric model, is commonly carried out via maximum likelihood (Olsson,, 1979). However, recent work has demonstrated that maximum likelihood (ML) estimation of polychoric correlation is highly sensitive to violations of the assumption of underlying normality. Violations of this assumption result in a misspecified polychoric model, which can lead to substantially biased estimates of its parameters and those of subsequent multivariate models, particularly SEMs (Grønneberg & Foldnes,, 2022; Foldnes & Grønneberg,, 2022, 2020, 2019; Jin & Yang-Wallentin,, 2017).
@@
To enhance accessibility and adoption by empirical researchers, an implementation of our proposed methodology in R (R Core Team,, 2024) is freely available in the package robcat (for “ROBust CATegorical data analysis”; Welz et al.,, 2025) on CRAN (the Comprehensive R Archive Network) at https://CRAN.R-project.org/package=robcat.
@@ 2 Related literature
Existing model-based approaches to account for careless responding in various models typically explicitly model carelessness trough mixture models (e.g., Steinmann et al.,, 2022; Van Laar & Braeken,, 2022; Ulitzsch et al., 2022a, ; Ulitzsch et al., 2022b, ; Arias et al.,, 2020). In contrast, our method does not model carelessness since we refrain from making assumptions on how the polychoric model might be misspecified. We refer to Alfons & Welz, (2024) for a detailed overview of methods addressing careless responding in various settings.
Conceptually related to our approach, Itaya & Hayashi, (2025) propose a way to robustly estimate parameters in item response theory (IRT) models. Their approach is conceptually similar to ours in the sense that it is based on minimizing a notion of divergence between an empirical density (from observed data) and a theoretical density of the IRT model. Like our approach, they achieve robustness by implicitly downweighting responses that the postulated model cannot fit well. Methodologically, our approach is different from Itaya & Hayashi, (2025) because our method is based on $C$-estimation (Welz,, 2024), which is designed specifically for categorical data, while they use density power divergence estimation (DPD) theory (Basu et al.,, 1998), which is not restricted to categorical data. [^2] Another slightly more subtle difference is that Itaya & Hayashi, (2025) robustify marginal maximum ML, whereas we robustify joint ML estimation. Marginal ML is often used to estimate IRT models because joint ML is known to yield inconsistent parameter estimates in such models (e.g., Lindsay et al.,, 1991). The polychoric correlation model does not suffer from such problems and can therefore be estimated via joint ML. A relevant consequence is that our estimator is fully efficient, whereas DPD estimators lose efficiency as a price for gaining robustness. To the best of our knowledge, DPD estimators have not yet been studied for estimating polychoric correlation.
@@ 3.1 The polychoric model
$X=\begin{cases}1&\text{if }\xi<a_{1},\\ 2&\text{if }a_{1}\leq\xi<a_{2},\\ 3&\text{if }a_{2}\leq\xi<a_{3},\\ \vdots\\ K_{X}&\text{if }a_{K_{X}-1}\leq\xi,\\ \end{cases}\qquad\textnormal{ and }\qquad Y=\begin{cases}1&\text{if }\eta<b_{1},\\ 2&\text{if }b_{1}\leq\eta<b_{2},\\ 3&\text{if }b_{2}\leq\eta<b_{3},\\ \vdots\\ K_{Y}&\text{if }b_{K_{Y}-1}\leq\eta,\\ \end{cases}$ (1)
@@
Combining the discretization model (1) with the latent normality model (2) yields the polychoric model. In this model, one refers to the correlation parameter $\rho=\mathrm{Cor}\left[\xi,\ \eta\right]$ as the polychoric correlation coefficient of the ordinal $X$ and $Y$. The polychoric model is subject to $d=K_{X}+K_{Y}-1$ parameters, namely the polychoric correlation coefficient from the latent normality model (2) and the two sets of thresholds from the discretization model (1). These parameters are jointly collected in a $d$-dimensional vector
@@
Regarding identification, it is worth mentioning that in the case where both $X$ and $Y$ are dichotomous, the polychoric model is exactly identified by the standard bivariate normal distribution. If at least one of the ordinal variables has more than two response categories, the polychoric model is over-identified, so it could identify more parameters than those in $\bm{\theta}$ [^3] An ordinal sample provides $K_{X}K_{Y}$ statistics, namely the observed frequencies for each response category $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Since the sum of the individual response frequencies must equal the sample size, one loses one degree of freedom so that only $K_{X}K_{Y}-1$ of the frequency statistics are independent. Subsequently, the ordinal sample can identify at most $K_{X}K_{Y}-1$ parameters. In particular, when $X$ and $Y$ are both dichotomous, only three parameters can be identified. In this case, the polychoric model (which reduces to the tetrachoric model here) depends on three parameters, namely a correlation parameter and two threshold parameters. We refer to Olsson, (1979, Section 2) for a related discussion.
@@ 3.2 Maximum likelihood estimation
[…]
@@ 4.1 Partial misspecification of the polychoric model
Although we make no assumption on the specific value of the contamination fraction $\varepsilon$ in the partial misspecification model (7), we require the identification restriction $\varepsilon\in[0,0.5)$. That is, we require that the polychoric model is correctly specified for the majority of observations, which is standard in the robust statistics literature (e.g., Hampel et al.,, 1986, p. 67). While it is in principle possible to also consider a contamination fraction between 0.50.50.5 and 111, one would need to impose certain additional assumptions on the correct model to distinguish it from incorrect ones when the majority of observations are not generated by the correct model. Since we prefer refraining from imposing additional assumptions, we only consider $\varepsilon\in[0,0.5)$. We discuss the link between identification and contamination fractions beyond 0.5 in more detail in Appendix E.1.
@@ 4.2 Response probabilities under partial misspecification
[…]
@@ 5.1 The estimator
is referred to as Pearson residual (PR) (Lindsay,, 1994). It takes values in [1,+)[-1,$\infty$)[ - 1 , + ∞ ) and can be interpreted as a goodness-of-fit measure. PR values close to 0 indicate a good fit between data and polychoric model at $\bm{\theta}$, whereas values toward 1-1- 1 or +$\infty$+ ∞ indicate a poor fit because empirical response probabilities disagree with their model counterparts. To achieve robustness to misspecification of the polychoric model, responses that cannot be modeled well by the polychoric model, as indicated by their PR being away from 0, should receive less weight in the estimation procedure such that they do not over-proportionally affect the fit. Downweighting when necessary is achieved through a specific choice of discrepancy function $\varphi(\cdot)$ proposed by Welz, (2024), which is a special case of a function suggested by Ruckstuhl & Welsh, (2001). The discrepancy function reads
@@ 5.2.2 Asymptotic analysis
as $N\to\infty$, where “$\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}$” and “$\stackrel{{\scriptstyle\mathrm{d}}}{{\longrightarrow}}$” denote convergence in probability and distribution, respectively. The asymptotic covariance matrix has a sandwich-type construction
@@
$\bm{M}\left(\bm{\theta}\right)=\frac{\partial^{2}L\big{(}\bm{\theta},f_{\varepsilon}\big{)}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}\quad\ \textnormal{and}\ \quad\bm{U}\left(\bm{\theta}\right)=\mathbb{V}\mathrm{ar}_{f_{\varepsilon}}\left[\frac{\partial\log\left(p_{XY}(\bm{\theta})\right)}{\partial\bm{\theta}}\mathds{1}\left\{\frac{f_{\varepsilon}\left(X,Y\right)}{p_{XY}}-1\in[-1,c]\right\}\right],$
respectively, are the Hessian matrix of the population loss and the covariance matrix (evaluated at $f_{\varepsilon}$) of the likelihood score function—that is, the gradient of $\log(p_{XY}(\bm{\theta}))$—weighted by stochastic binary weights whether the Pearson residual is smaller than or equal to the tuning constant $c$ [^9] It is worth mentioning that for the MLE ($c=+\infty$), our sandwich-type covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ reduces to that of White, (1982) and Huber, (1967), who studied the limit distribution of ML in misspecified models. We derive closed-form expressions of the matrices $\bm{M}\left(\bm{\theta}\right)$ and $\bm{U}\left(\bm{\theta}\right)$ as well as their properties in Appendix A.
@@ 6 Simulation studies on partial misspecification
In this section, we employ two simulation studies to demonstrate the robustness gain of our proposed estimator under partial misspecification of the polychoric model. The first simulation design (Section 6.1) is a simplified setting with respect to the partial misspecification, chosen specifically to illustrate the effects of a particular type of contamination with high leverage affecting only a small number of contingency table cells. The second design (Section 6.2) is more involved and considers estimation of a polychoric correlation matrix with a contamination type that scatters in many directions so that nonnormal data points can occur in every contingency table cell. Section 6.3 summarizes findings from additional simulations in the appendix. For all simulation designs, we perform 5,0005,0005 , 000 repetitions.
@@ 6.1 Individual polychoric correlation coefficient
[…]
@@
Figure 4: Boxplot visualization of the bias of three estimators of the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{*}$, for various contamination fractions in the misspecified polychoric model across 5,000 repetitions. The estimators are the robust estimator with $c=0.6$ (left), the MLE (center), and the Pearson sample correlation (right). Diamonds represent the respective average bias. The dashed line denotes value 0 and the dotted line $-\rho_{*}=-0.5$, the latter of which indicating a sign flip in the correlation estimate.
@@
$\varepsilon=0$ Robust 0.504 0.004 0.027 0.027 0.001-0.001- 0.001 0.931 0.104
ML 0.500 0.000 0.027 0.026 0.001-0.001- 0.001 0.939 0.102
Sample cor 0.456 0.044-0.044- 0.044 0.025 0.025 0.003 0.690 0.110
$\varepsilon=0.01$ Robust 0.501 0.001 0.028 0.027 0.001-0.001- 0.001 0.938 0.107
ML 0.441 0.059-0.059- 0.059 0.027 0.028 0.001 0.446 0.110
Sample cor 0.402 0.098-0.098- 0.098 0.025 0.029 0.004 0.043 0.114
$\varepsilon=0.05$ Robust 0.495 0.005-0.005- 0.005 0.029 0.029 0.001-0.001- 0.001 0.939 0.112
ML 0.252 0.248-0.248- 0.248 0.028 0.033 0.005 0.000 0.128
Sample cor 0.221 0.279-0.279- 0.279 0.025 0.031 0.006 0.000 0.121
$\varepsilon=0.1$ Robust 0.488 0.012-0.012- 0.012 0.031 0.031 0.000 0.933 0.120
ML 0.062 0.438-0.438- 0.438 0.028 0.035 0.006 0.000 0.135
Sample cor 0.052 0.448-0.448- 0.448 0.025 0.032 0.007 0.000 0.124
$\varepsilon=0.15$ Robust 0.480 0.020-0.020- 0.020 0.033 0.033 0.000 0.907 0.130
ML 0.097-0.097- 0.097 0.597-0.597- 0.597 0.027 0.035 0.007 0.000 0.135
Sample cor 0.076-0.076- 0.076 0.576-0.576- 0.576 0.025 0.032 0.007 0.000 0.124
$\varepsilon=0.2$ Robust 0.472 0.028-0.028- 0.028 0.036 0.036 0.001 0.882 0.142
ML 0.229-0.229- 0.229 0.729-0.729- 0.729 0.026 0.033 0.008 0.000 0.131
Sample cor 0.176-0.176- 0.176 0.676-0.676- 0.676 0.025 0.031 0.007 0.000 0.122
$\varepsilon=0.3$ Robust 0.450 0.050-0.050- 0.050 0.043 0.045 0.002 0.789 0.176
ML 0.431-0.431- 0.431 0.931-0.931- 0.931 0.022 0.030 0.008 0.000 0.118
Sample cor 0.322-0.322- 0.322 0.822-0.822- 0.822 0.024 0.030 0.007 0.000 0.117
$\varepsilon=0.4$ Robust 0.413 0.087-0.087- 0.087 0.053 0.067 0.014 0.595 0.261
ML 0.571-0.571- 0.571 1.071-1.071- 1.071 0.019 0.026 0.008 0.000 0.103
Sample cor 0.424-0.424- 0.424 0.924-0.924- 0.924 0.024 0.029 0.005 0.000 0.112
$\varepsilon=0.49$ Robust 0.357 0.143-0.143- 0.143 0.062 0.071 0.009 0.247 0.278
ML 0.662-0.662- 0.662 1.162-1.162- 1.162 0.016 0.024 0.008 0.000 0.095
Sample cor 0.492-0.492- 0.492 0.992-0.992- 0.992 0.023 0.028 0.004 0.000 0.108
@@
Table 1: Results for the robust estimator with $c=0.6$, the MLE, and the Pearson sample correlation, for various contamination fractions across 5,000 simulated datasets. The true polychoric correlation coefficient is $\rho_{*}=0.5$. We compute the average of point estimates $\widehat{\rho}_{N}$ of the polychoric correlation coefficient, the average bias ($\widehat{\rho}_{N}-\rho_{*}$), the standard deviation of the $\widehat{\rho}_{N}$ (SE; an approximation of the true standard error), the average standard error estimate $\widehat{\text{SE}}$, the corresponding average bias ($-\text{SE}$ SE), confidence interval coverage with respect to the true $\rho_{*}$ at nominal level 95%, and the average length of the respective confidence intervals.
@@
Figure 4 visualizes the bias of each estimator with respect to the true polychoric correlation $\rho_{*}$ across the 5,000 simulated datasets. An analogous plot for the whole parameter $\bm{\theta}_{*}$ can be found in Appendix D.1; the results are similar to those of $\rho_{*}$. Additional performance measures are shown in Table 1. For all considered contamination fractions, the estimates of the MLE and sample correlation are somewhat similar, which is expected because these two estimators are known to yield similar results when there are five or more rating options and the discretization thresholds are symmetric (cf., Rhemtulla et al.,, 2012). In the absence of contamination, the MLE and the robust estimator yield accurate estimates. Both estimators are nearly equivalent to one another in the sense that their point estimates, standard deviation, and coverage at significance level $\alpha=0.05$ are very similar. However, when contamination is introduced, MLE, sample correlation, and the robust estimator yield noticeably different results. Already at the small contamination fraction $\varepsilon=0.01$ (corresponding to only 10 observations), MLE and sample correlation are noticeably biased, resulting in poor coverage of only about 0.45 and 0.04, respectively. Increasing the contamination fraction to the still relatively small value of $\varepsilon=0.05$, MLE and sample correlation start to be substantially biased with average biases of 0.25-0.25- 0.25 and 0.28-0.28- 0.28, respectively, leading to zero coverage. The biases of these two methods further deteriorate as the contamination fraction is gradually increased. At $\varepsilon\geq 0.15$, MLE and sample correlation produce estimates that are not only severely biased but also sign-flipped: while the true correlation is positive (0.5), both estimates are negative. In stark contrast, the proposed robust estimator remains accurate throughout nearly all considered contamination fractions. At the small $\varepsilon=0.01$, the robust estimator is nearly unaffected, while at $\varepsilon=0.15$, it only exhibits a minor bias of about 0.02-0.02- 0.02. Even at extreme contamination $\varepsilon=0.4$, its bias amounts to less than 0.1. In addition, coverage of the robust method remains above or close to 0.9 for contamination fractions $\varepsilon\leq 0.2$.

### v5

@@ 1 Introduction
Ordinal data are ubiquitous in psychology and related fields. With such data, e.g., arising from responses to rating scales, it is often recommended to estimate correlation matrices through polychoric correlation coefficients (e.g., Foldnes & Grønneberg,, 2022; Garrido et al.,, 2013; Holgado-Tello et al.,, 2010). The resulting polychoric correlation matrix is an important building block in subsequent multivariate models like factor analysis models and structural equation models (SEMs), as well as in exploratory methods like principal component analysis, multidimensional scaling, and clustering techniques (see, e.g., Mair,, 2018, for an overview). An individual polychoric correlation coefficient is the population correlation between two underlying latent variables that are postulated to have generated the observed categorical data through an unobserved discretization process. Traditionally, it is assumed that the two latent variables are standard bivariate normally distributed (Pearson,, 1901) to estimate the polychoric correlation coefficient from observed ordinal data. Estimation of this latent normality model, called the polychoric model, is commonly carried out via maximum likelihood (Olsson,, 1979). However, recent work has demonstrated that maximum likelihood (ML) estimation of polychoric correlation is highly sensitive to violations of the assumption of underlying normality. Violations of this assumption result in a misspecified polychoric model, which can lead to substantially biased estimates of its parameters and those of subsequent multivariate models (Grønneberg & Foldnes,, 2022; Foldnes & Grønneberg,, 2020, 2019; Jin & Yang-Wallentin,, 2017), particularly SEMs using diagonally weighted least squares (Foldnes & Grønneberg,, 2022), where the latter is based on weights derived under latent normality.
@@
To enhance accessibility and adoption by empirical researchers, an implementation of our proposed methodology in R (R Core Team,, 2024) is freely available in the package robcat (for “ROBust CATegorical data analysis”; Welz et al.,, 2025) on CRAN (the Comprehensive R Archive Network) at https://CRAN.R-project.org/package=robcat. Replication materials for all numerical results in this paper are provided on GitHub at https://github.com/mwelz/robust-polycor-replication.
@@ 2 Related literature
Existing model-based approaches to account for careless responding in various models typically explicitly model carelessness trough mixture models (e.g., Steinmann et al.,, 2022; Van Laar & Braeken,, 2022; Ulitzsch et al., 2022a, ; Ulitzsch et al., 2022b, ; Arias et al.,, 2020). In contrast, our method does not model carelessness since we refrain from making assumptions on how the polychoric model might be misspecified. Another way to address careless responding is to directly detect them through person-fit indices and subsequently remove them from the sample (e.g., Patton et al.,, 2019). As primary difference, our method simultaneously downweights aberrant observations during estimation rather than removing them. We refer to Alfons & Welz, (2024) for a detailed overview of methods addressing careless responding in various settings.
Conceptually related to our approach, Itaya & Hayashi, (2025) propose a way to robustly estimate parameters in item response theory (IRT) models. Their approach is conceptually similar to ours in the sense that it is based on minimizing a notion of divergence between an empirical density (from observed data) and a theoretical density of the IRT model. Like our approach, they achieve robustness by implicitly downweighting responses that the postulated model cannot fit well. Methodologically, our approach is different from Itaya & Hayashi, (2025) because our method is based on $C$-estimation (Welz,, 2024), which is designed specifically for categorical data, while they use density power divergence estimation (DPD) theory (Basu et al.,, 1998), which is not restricted to categorical data. [^2] Another slightly more subtle difference is that Itaya & Hayashi, (2025) robustify marginal maximum ML, whereas we robustify joint ML estimation. Marginal ML is often used to estimate IRT models because joint ML is known to yield inconsistent parameter estimates in such models (e.g., Lindsay et al.,, 1991). The polychoric correlation model does not suffer from such problems and can therefore be estimated via joint ML. Furthermore, marginal ML in IRT is a full information estimation concept, whereas multivariate models that are being fitted to a given polychoric correlation matrix require limited information estimation. Hence, the work of Itaya & Hayashi, (2025) demonstrates that model misspecification is also a relevant issue in full information estimation. A relevant consequence is that our estimator is fully efficient, whereas DPD estimators lose efficiency as a price for gaining robustness. To the best of our knowledge, DPD estimators have not yet been studied for estimating polychoric correlation.
@@ 3.1 The polychoric model
[…]
@@
Combining the discretization model (1) with the latent normality model (2) yields the polychoric model. In this model, one refers to the correlation parameter $\rho=\mathbb{C}\mathrm{or}\left[\xi,\ \eta\right]$ as the polychoric correlation coefficient of the ordinal $X$ and $Y$. The polychoric model is subject to $d=K_{X}+K_{Y}-1$ parameters, namely the polychoric correlation coefficient from the latent normality model (2) and the two sets of thresholds from the discretization model (1). These parameters are jointly collected in a $d$-dimensional vector
@@
Regarding identification, it is worth mentioning that in the case where both $X$ and $Y$ are dichotomous, the polychoric model is exactly identified by the standard bivariate normal distribution. If at least one of the ordinal variables has more than two response categories, the polychoric model is over-identified, so it could identify more parameters than those in $\bm{\theta}$. [^3] An ordinal sample provides $K_{X}K_{Y}$ statistics, namely the observed frequencies for each response category $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Since the sum of the individual response frequencies must equal the sample size, one loses one degree of freedom so that only $K_{X}K_{Y}-1$ of the frequency statistics are independent. Subsequently, the ordinal sample can identify at most $K_{X}K_{Y}-1$ parameters. In particular, when $X$ and $Y$ are both dichotomous, only three parameters can be identified. In this case, the polychoric model (which reduces to the tetrachoric model here) depends on three parameters, namely a correlation parameter and two threshold parameters. We refer to Olsson, (1979, Section 2) for a related discussion.
@@ 3.2 Maximum likelihood estimation
Suppose we observe a sample $\{(X_{i},Y_{i})\}_{i=1}^{N}$ of $N$ independent copies of $(X,Y)$ generated by the polychoric model under the true parameter $\bm{\theta}_{*}$. The sample may be observed directly or as a $K_{X}\times K_{Y}$ contingency table that cross-tabulates the observed frequencies. Denote by
@@
Software implementations of polychoric correlation vary with respect to their estimation strategy. For instance, the popular R packages lavaan (Rosseel,, 2012) and psych (Revelle,, 2024) only support the 2-step approach, while the package polycor (Fox,, 2022) supports both the 2-step-approach and simultaneous estimation of all model parameters, with the former being the default. Our implementation of ML estimation in package robcat also supports both strategies.
@@ 4.1 Partial misspecification of the polychoric model
Although we make no assumption on the specific value of the contamination fraction $\varepsilon$ in the partial misspecification model (7), we require the identification restriction $\varepsilon\in[0,0.5)$. That is, we require that the polychoric model is correctly specified for the majority of observations, which is standard in the robust statistics literature (e.g., Hampel et al.,, 1986, p. 67). While it is in principle possible to also consider a contamination fraction between $0.5$ and $1$, one would need to impose certain additional assumptions on the correct model to distinguish it from incorrect ones when the majority of observations are not generated by the correct model. Since we prefer refraining from imposing additional assumptions, we only consider $\varepsilon\in[0,0.5)$. We discuss the link between identification and contamination fractions beyond 0.5 in more detail in Appendix E.1.
@@ 4.2 Response probabilities under partial misspecification
Figure 1 visualizes a simulated example of bivariate data drawn from contaminated distribution $G_{\varepsilon}$, where a fraction of $\varepsilon=0.15$ of the data follow a bivariate normal contamination distribution $H$ (orange dots) with mean $(2.5,-2.5)^{\top}$, variance $(0.25,0.25)^{\top}$, and zero correlation, whereas the remaining data are generated by a standard bivariate normal distribution with correlation $\rho_{*}=0.5$ (gray dots). In this example, the data from the contamination distribution $H$ primarily inflate the cell $(x,y)=(5,1)$ after discretization. That is, this cell will have a larger empirical frequency than the polychoric model allows for, since the probability of this cell is nearly zero at the polychoric model, yet many realized responses will populate it. Consequently, due to (partial) misspecification of the polychoric model, a maximum likelihood estimate of $\rho_{*}$ on these data might be substantially biased for $\rho_{*}$. Indeed, calculating the MLE using the data plotted in Figure 1 yields an estimate of $\widehat{\rho}_{N}^{\mathrm{\ MLE}}=-0.10$, which is far off from the true $\rho_{*}=0.5$. In contrast, our proposed robust estimator, which is calculated from the exact same information as the MLE and is defined in Section 5, yields a fairly accurate estimate of $0.47$.
@@ 5.1 The estimator
is referred to as Pearson residual (PR) (Lindsay,, 1994). It takes values in $1,+\infty)$ and can be interpreted as a goodness-of-fit measure. PR values close to 0 indicate a good fit between data and polychoric model at $\bm{\theta}$, whereas values toward 1-1 or $\infty$ indicate a poor fit because empirical response probabilities disagree with their model counterparts. To achieve robustness to misspecification of the polychoric model, responses that cannot be modeled well by the polychoric model, as indicated by their PR being away from 0, should receive less weight in the estimation procedure such that they do not over-proportionally affect the fit. Downweighting when necessary is achieved through a specific choice of discrepancy function $\varphi(\cdot)$ proposed by Welz, (2024), which is a special case of a function suggested by Ruckstuhl & Welsh, (2001). The discrepancy function reads
@@ 5.2.2 Asymptotic analysis
[…]
@@
$\bm{M}\left(\bm{\theta}\right)=\frac{\partial^{2}L\big(\bm{\theta},f_{\varepsilon}\big)}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}\quad\ \textnormal{and}\ \quad\bm{U}\left(\bm{\theta}\right)=\mathbb{V}\mathrm{ar}_{f_{\varepsilon}}\left[\frac{\partial\log\left(p_{XY}(\bm{\theta})\right)}{\partial\bm{\theta}}\mathds{1}\left\{\frac{f_{\varepsilon}\left(X,Y\right)}{p_{XY}(\bm{\theta})}-1\in[-1,c]\right\}\right],$
respectively, are the Hessian matrix of the population loss and the covariance matrix (evaluated at $f_{\varepsilon}$) of the likelihood score function—that is, the gradient of $\log(p_{XY}(\bm{\theta}))$—weighted by stochastic binary weights whether the Pearson residual is smaller than or equal to the tuning constant cc. [^9] It is worth mentioning that for the MLE ($c=+\infty$), our sandwich-type covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ reduces to that of White, (1982) and Huber, (1967), who studied the limit distribution of ML in misspecified models. We derive closed-form expressions of the matrices $\bm{M}\left(\bm{\theta}\right)$ and $\bm{U}\left(\bm{\theta}\right)$ as well as their properties in Appendix A.
@@ 6 Simulation studies on partial misspecification
In this section, we employ two simulation studies to demonstrate the robustness gain of our proposed estimator under partial misspecification of the polychoric model. The first simulation design (Section 6.1) is a simplified setting with respect to the partial misspecification, chosen specifically to illustrate the effects of a particular type of contamination with high leverage affecting only a small number of contingency table cells. The second design (Section 6.2) is more involved and considers estimation of a polychoric correlation matrix with a contamination type that scatters in many directions so that nonnormal data points can occur in every contingency table cell. Section 6.3 summarizes findings from additional simulations in the appendix. For all simulation designs, we perform $5,000$ repetitions.
@@ 6.1 Individual polychoric correlation coefficient
and let the true polychoric correlation coefficient in latent normality model (2) be $\rho_{*}=0.5$. To simulate partial misspecification of the polychoric model according to (7), we let a fraction $\varepsilon$ of the data be generated by a particular contamination distribution HH—which is left unspecified and therefore not explicitly modeled by our estimator—namely a bivariate normal distribution with mean $(2.5,-2.5)^{\top}$, variances $(0.25,0.25)^{\top}$, and zero covariance (and therefore zero correlation). We discretize the realizations of the contamination distribution according to the same thresholds $a_{*,1},\dots,a_{*,4},b_{*,1},\dots,b_{*,4}$ as the uncontaminated realizations. This contamination distribution will inflate the empirical frequency of contingency table cells $(x,y)\in\{(5,1),(4,1),(5,2)\}$, in the sense that they have a higher realization probability than under the true polychoric model. [^11] Conceptually, contamination in the cell $(5,1)$ would correspond to a straightlining careless respondent in $(5,5)$ or $(1,1)$ if one of the items is negatively keyed. Thus, after recoding, cell $(5,1)$ will be inflated. In fact, the data plotted in Figure 1 were generated by this process for contamination fraction $\varepsilon=0.15$, and one can see in this figure that particularly cell $(x,y)=(5,1)$ is sampled frequently although it only has a near-zero probability at the true polychoric model. The data points causing these three cells to be inflated are instances of negative leverage points. Here, such leverage points drag correlational estimates away from a positive value towards zero or, if there are sufficiently many of them, even a negative value. For intuition, one may think of such points as the responses of careless or inattentive participants whose responses are not based on item content. Although careless responding is only one special case of the unlimited and unrestricted variety of uninformative responses generated by $H$, we use careless responding as an illustrative running example throughout our simulations.
@@
Figure 4: Boxplot visualization of the bias of three estimators of the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{*}$, for various contamination fractions in the misspecified polychoric model across 5,000 repetitions. The estimators are the robust estimator with $c=0.6$ (left), the MLE (center), and the Pearson sample correlation (right). Diamonds represent the respective average bias. The dashed line denotes value 0 and the dotted line $\rho_{*}=-0.5$, the latter of which indicating a sign flip in the correlation estimate.
@@
$\varepsilon=0$ Robust 0.504 0.004 0.027 0.027 0.001-0.001 0.931 0.104
ML 0.500 0.000 0.027 0.026 0.001-0.001 0.939 0.102
Sample cor 0.456 0.044-0.044 0.025 0.025 0.003 0.690 0.110
$\varepsilon=0.01$ Robust 0.501 0.001 0.028 0.027 0.001-0.001 0.938 0.107
ML 0.441 0.059-0.059 0.027 0.028 0.001 0.446 0.110
Sample cor 0.402 0.098-0.098 0.025 0.029 0.004 0.043 0.114
$\varepsilon=0.05$ Robust 0.495 0.005-0.005 0.029 0.029 0.001-0.001 0.939 0.112
ML 0.252 0.248-0.248 0.028 0.033 0.005 0.000 0.128
Sample cor 0.221 0.279-0.279 0.025 0.031 0.006 0.000 0.121
$\varepsilon=0.1$ Robust 0.488 0.012-0.012 0.031 0.031 0.000 0.933 0.120
ML 0.062 0.438-0.438 0.028 0.035 0.006 0.000 0.135
Sample cor 0.052 0.448-0.448 0.025 0.032 0.007 0.000 0.124
$\varepsilon=0.15$ Robust 0.480 0.020-0.020 0.033 0.033 0.000 0.907 0.130
ML 0.097-0.097 0.597-0.597 0.027 0.035 0.007 0.000 0.135
Sample cor 0.076-0.076 0.576-0.576 0.025 0.032 0.007 0.000 0.124
$\varepsilon=0.2$ Robust 0.472 0.028-0.028 0.036 0.036 0.001 0.882 0.142
ML 0.229-0.229 0.729-0.729 0.026 0.033 0.008 0.000 0.131
Sample cor 0.176-0.176 0.676-0.676 0.025 0.031 0.007 0.000 0.122
$\varepsilon=0.3$ Robust 0.450 0.050-0.050 0.043 0.045 0.002 0.789 0.176
ML 0.431-0.431 0.931-0.931 0.022 0.030 0.008 0.000 0.118
Sample cor 0.322-0.322 0.822-0.822 0.024 0.030 0.007 0.000 0.117
$\varepsilon=0.4$ Robust 0.413 0.087-0.087 0.053 0.067 0.014 0.595 0.261
ML 0.571-0.571 1.071-1.071 0.019 0.026 0.008 0.000 0.103
Sample cor 0.424-0.424 0.924-0.924 0.024 0.029 0.005 0.000 0.112
$\varepsilon=0.49$ Robust 0.357 0.143-0.143 0.062 0.071 0.009 0.247 0.278
ML 0.662-0.662 1.162-1.162 0.016 0.024 0.008 0.000 0.095
Sample cor 0.492-0.492 0.992-0.992 0.023 0.028 0.004 0.000 0.108
@@
Table 1: Results for the robust estimator with $c=0.6$, the MLE, and the Pearson sample correlation, for various contamination fractions across 5,000 simulated datasets. The true polychoric correlation coefficient is $\rho_{*}=0.5$. We compute the average of point estimates $\widehat{\rho}_{N}$ of the polychoric correlation coefficient, the average bias ($\widehat{\rho}_{N}-\rho_{*}$), the standard deviation of the $\widehat{\rho}_{N}$ (SE; an approximation of the true standard error), the average standard error estimate $\widehat{\text{SE}}$, the corresponding average bias ($\widehat{\text{SE}}-\text{SE}$), confidence interval coverage with respect to the true $\rho_{*}$ at nominal level 95%, and the average length of the respective confidence intervals.
@@
Figure 4 visualizes the bias of each estimator with respect to the true polychoric correlation $\rho_{*}$ across the 5,000 simulated datasets. An analogous plot for the whole parameter $\bm{\theta}_{*}$ can be found in Appendix D.1; the results are similar to those of $\rho_{*}$. Additional performance measures are shown in Table 1. For all considered contamination fractions, the estimates of the MLE and sample correlation are somewhat similar, which is expected because these two estimators are known to yield similar results when there are five or more rating options and the discretization thresholds are symmetric (cf., Rhemtulla et al.,, 2012). In the absence of contamination, the MLE and the robust estimator yield accurate estimates. Both estimators are nearly equivalent to one another in the sense that their point estimates, standard deviation, and coverage at significance level $\alpha=0.05$ are very similar. However, when contamination is introduced, MLE, sample correlation, and the robust estimator yield noticeably different results. Already at the small contamination fraction $\varepsilon=0.01$ (corresponding to only 10 observations), MLE and sample correlation are noticeably biased, resulting in poor coverage of only about 0.45 and 0.04, respectively. Increasing the contamination fraction to the still relatively small value of $\varepsilon=0.05$, MLE and sample correlation start to be substantially biased with average biases of 0.25-0.25 and 0.28-0.28, respectively, leading to zero coverage. The biases of these two methods further deteriorate as the contamination fraction is gradually increased. At $\varepsilon\geq 0.15$, MLE and sample correlation produce estimates that are not only severely biased but also sign-flipped: while the true correlation is positive (0.5), both estimates are negative. In stark contrast, the proposed robust estimator remains accurate throughout nearly all considered contamination fractions. At the small $\varepsilon=0.01$, the robust estimator is nearly unaffected, while at $\varepsilon=0.15$, it only exhibits a minor bias of about 0.02-0.02. Even at extreme contamination $\varepsilon=0.4$, its bias amounts to less than 0.1. In addition, coverage of the robust method remains above or close to 0.9 for contamination fractions $\varepsilon\leq 0.2$.